Trace the curve Y=(2x²-8)/(x²-16) by showing all the properties used to trace the curve.
1
Expert's answer
2020-02-27T11:52:50-0500

1) Domain, vertical asymptotes and intercepts.

The domain is "(-\infty,-4) \cup (-4, 4) \cup (4, \infty)," since "x" cannot be equal to "\pm 4."

Determine the one-sided limits in points "-4" and "4."


"\lim\limits_{x\to -4^{+}}\frac{2x^2-8}{x^2-16} = -\infty \quad
\lim\limits_{x\to -4^{-}}\frac{2x^2-8}{x^2-16} = +\infty \\

\lim\limits_{x\to 4^{+}}\frac{2x^2-8}{x^2-16} = +\infty \quad
\lim\limits_{x\to 4^{-}}\frac{2x^2-8}{x^2-16} = -\infty"


Therefore, "x=-4 \text{ and } x=4" are the vertical asymptotes.

Determine the "x\text{-intercept(s)}."

"y=0 \\
\frac{2x^2-8}{x^2-16} = 0 \\
2x^2 - 8 = 0 \\
2(x^2-4)=0 \\
(x-2)(x+2) = 0 \\
x=-2, \;x=2"

So, the "x\text{-intercepts are } (-2,0) \text{ and } (2,0)."


Determine the "y\text{-intercept(s)}."

"x=0 \\
y = \frac{2(0)^2 - 8}{(0)^2-16} = \frac{1}{2}"

So, the "y\text{-intercept is } (0,\frac{1}{2})."


2) Symmetry.

"f(-x) = \frac{2(-x)^2-8}{(-x)^2-16} = \frac{2x^2-8}{x^2-16} = f(x)"


Therefore, the function is even.


3) Non-vertical asymptotes.

Let "y=kx+b" is an equation of a non-vertical asymptote. Here,


"k = \lim\limits_{x\to\infty}\frac{f(x)}{x} = \lim\limits_{x\to\infty}\frac{2x^2-8}{x(x^2-16)} = 0 \\

b = \lim\limits_{x\to\infty}(f(x) - kx) = \lim\limits_{x\to\infty}\frac{2x^2-8}{x^2-16} = 2"

Therefore, "y=2" is the horizontal asymptote.


4) Monotonicity

"y'=(\frac{2x^2-8}{x^2-16})' = \frac{4x(x^2-16) - 2x(2x^2-8)}{(x^2-16)^2} = \frac{-48x}{(x^2-16)^2} \\

y'=0 \text{ when } x=0"

If "x<0", then "y'>0". If "x>0", then "y'<0." Therefore, when "x<0" the function is increasing, when "x>0" the function is decreasing, and "x=0" is the maximum.


5) Convex

"y'' = (\frac{2x^2-8}{x^2-16})'' = (\frac{-48x}{(x^2-16)^2})' = \frac{-48(x^2-16)^2 + 48x\cdot2(x^2-16)\cdot(2x)}{(x^2-16)^4} \\

\quad = \frac{48(3x^2+16)}{(x^2-16)^3}"

The second derivative has no roots, but it has the points, when the second derivative does not exist "(x=-4,x=4)."


If "x<-4", then "y''>0." If "-4<x<4", then "y''<0." If "x>4," then "y''>0." Therefore, when "x<-4 \text{ and } x>4" the function concave upward, and when "-4<x<4" the function concave downward.


Draw a graph.